There is 24 feet of fabric to be cut to make curtains. If every curtain is going to be 2 and 2/3 feet, how many curtains can be made?
1
Expert's answer
2012-08-16T07:27:56-0400
To calculate the amount of curtains we should divide the whole amount of fabric by the length of one curtain:
N=24/(2 2/3)=24/(8/3)=24*3/8=9